OpenStudy (anonymous):

Algebra Urgent: The sum of the ages of Alex and Harry together is 36. If Alex's current age is 3 times as much as Harry's age was 8 years ago, how old is each?

OpenStudy (unklerhaukus):

A+H=36 A=3(H-8)

OpenStudy (goformit100):

let harry age = x then alex age is 3(x-8) therefore a+x=36 a=3(x-8)

OpenStudy (anonymous):

Thanks you both, but the thing is i know that part, i need help on how to get the solution like Alex's and Harry's age each.

OpenStudy (unklerhaukus):

substitute the second equation into the firs equation {3(H-8)}+H=36 now you can solve for h ,

OpenStudy (anonymous):

Oh, thanks!!!

OpenStudy (unklerhaukus):

{3(H-8)}+H=36 3H-24+H=36 4H=36+24 4H=60 H=15 now plug H=15 into the first equation to solve for A

OpenStudy (anonymous):

Yeah i got it :) Thank you so much xD

OpenStudy (unklerhaukus):

how old did you determine alex to be/

OpenStudy (anonymous):

21

OpenStudy (anonymous):

so 15+21=36 it works :)

OpenStudy (unklerhaukus):

A=3(H-8) 21=3(15-8)=3x7 too you have found the correct solution

OpenStudy (anonymous):

Tysm :)